# Jon Danielsson London School of Economics Andreas Uthemann

### 10 Jan 2024 v3

@@ 1 Introduction
Artificial intelligence (AI) will transform how the financial system is governed, promising improved efficiency, robustness and impartiality at much lower costs than existing arrangements. AI also threatens financial stability because of data limitations, misspecified objectives, unpredictable behaviour when confronted with unknown-unknowns, misaligned incentives and complex strategic interactions with humans and other AIs.
A financial authority entrusted with regulating the financial system has two objectives. The first is microprudential regulation, or micropru, concerned with day-to-day issues such as risk management, consumer protection and fraud. AI will benefit the micropru authorities as data is ample, rules are mostly fixed on the timescale decisions are made, and the cost of mistakes is small.
The second policy objective is macroprudential regulations, or macropru, focusing on broad-picture issues that relate to financial stability. The emphasis is decidedly long run, to avoid large losses and systemic financial crises years or even decades into the future and to resolve such crises when they occur. Macropru is more difficult to execute and less accurate than micropru, as the events under control are infrequent and often unique in essential aspects, while the financial markets and regulations are continually evolving. That implies considerable uncertainty about the ability of AI in this problem domain and the dangers arising from its use, as noted by Danielsson, Macrae and Uthemann (2022).
The usefulness of AI for any task depends on the structure of the task at hand, and it is helpful to think of AI’s application to financial policy on a spectrum. At one end, we have a problem with ample data and clear and simple objectives, playing to the strength of AI and its learning algorithms. As the frequency of events drops and rules are increasingly mutable, the AI’s advantages erode until we reach once-in-a-working-lifetime events, such as having to resolve a systemic financial crisis, where AI, though potentially very valuable, can threaten the system’s stability if used inappropriately.
That suggests six questions to ask when evaluating the use of AI for financial regulation and crisis resolution. Table 1 on page 7 shows how each relates to particular activities, ranging from routine risk management to systemic crises.
  1. 1.
    Does the AI have enough data?
  2. 2.
    Are the rules of the problem domain immutable?
  3. 3.
    Can AI be given clear objectives?
  4. 4.
    Does the authority the AI works for make decisions on its own?
  5. 5.
    Can we attribute responsibility for misbehaviour and mistakes?
  6. 6.
    Are the consequences of mistakes catastrophic?
@@
There is no single notion of AI, but it is generally seen as a computer algorithm that can perform tasks that humans typically do. What distinguishes AI from machine learning (ML) and traditional statistics is that it not merely provides quantitative inputs for human analysis but also gives recommendations and makes decisions. [^1] For general overview of AI Russel (2019) is particularly useful. For a more technical review of the underlying ML techniques, see Murphy (2023) and Prince (2023). Today’s AI excel at extracting patterns from large, unstructured datasets and use these for classification and prediction tasks. But they do not reason in the same way humans do.
Once AI achieves proficiency at a given task — has been trained — it can use its trained model to advise human decision makers or make decisions autonomously. As AI meeting objectives given its understanding of the environment and the economic notion of agents maximising expected utility given beliefs are conceptually closely related, it is particularly beneficial to rely on Russel and Norvig’s (2010) notion of AI as a rational maximising agent in the analysis of AI use in financial policy and we adopt this perspective in what follows.
@@
Private sector financial institutions are increasingly using AI in risk management, credit allocation, fraud detection and regulatory compliance. Even if the more traditional banks have been reticent in adopting AI, some banks employ large AI teams, and challenger banks have been particularly enthusiastic about its use. Significant cost savings and efficiency improvements in a highly competitive landscape will likely rapidly drive the further adoption of AI in the private sector.
@@
The financial authorities need to respond, and most are formulating AI policy (see, e.g. Moufakkir, 2023). At the same time, some authorities might prefer a slow, deliberative and conservative approach that will not be tenable for three reasons. First, an authority that persists in regulating with traditional methods may find itself outmanoeuvred. Second, a policy of not using AI for high-level decisions will likely be undermined by AI being adopted by stealth. Finally, AI may be considered essential for advising on a critical task. Even if an authority wanted to limit AI to advisory roles, there might not be much difference between AI making decisions and AI providing crucial advice, particularly if AI’s internal representation of the system is not intelligible to its human operators.
@@
The usefulness of AI for the financial authorities is directly affected by five conceptual challenges: The first, perhaps paradoxically, is data. After all, the financial system may appear to be the ideal use case for AI because it creates vast amounts of data, leaving plenty for AI to train on. However, data collection is incomplete, and the recorded data are often inconsistently or even inaccurately measured, leading to non-representative samples. What is more, financial crises are rare, meaning there is little data available on the most important events for macropru.
@@
The second challenge arises from the uniqueness of crises. While major financial crises share common fundamental features, every crisis is unique because the driving factors are specific to the institutional structure of that time and the particular regulatory and political regimes in place. Rarity and essential uniqueness in a sparsely monitored system imply that systemic crises can be seen as “uncertain” in Frank Knight’s (1921) classification and thus do not fit well into the conventional ML paradigm.
@@
AI’s third conceptual challenge relates to how the financial system responds to control, echoing Goodhart’s (1974) law and the Lucas (1976) critique. New regulations change the structure of the financial system, making it difficult for AI to learn from historical data about the possible consequences of regulatory actions.
@@
The fourth conceptual problem stems from a lack of clarity on what objectives the AI should pursue. This is particularly problematic for macropru, where the regulators’ problem is often described in abstract high-level terms such as “ensure financial stability”.
The final conceptual challenge relates to difficulties aligning AI’s incentives with the authorities’ objectives.
@@
The literature on AI has identified particular risks that AI engines pose; see, for example, Barnichon et al. (2022), and several researchers, such as Weidinger et al. (2022), Bengio et al. (2023) and Shevlane et al. (2023) identify societal risks arising from AI. When considering how these interact with the conceptual challenges listed above, we find five channels for how AI use may destabilise the financial system.
@@
The first is the malicious use of AI, which is a particular concern in the financial system because it is replete with highly resourced profit maximising economic agents. They can change the system in a way that benefits them while being difficult to detect by competitors and regulators. They may even seek to deliberately create market stress. Malicious agents might either directly manipulate AI engines or use them to find loopholes to evade control. Such activities can be socially undesirable and even run against the interests of the institution operating the AI engine. Apart from malicious uses by rogue traders and criminals for profit purposes, we also have contemd with terrorists and nation-states seeking to create social disorder.
The second channel is misinformed use and an overreliance on AI, a particular problem if AI is used when objectives are unclear and data-driven algorithms are at risk of incorrectly extrapolating because of insufficient data. Technology is often initially met with scepticism, but as it comes to be seen as performing better than what came before, it is increasingly trusted. AI earns trust by successfully executing those tasks best suited for it, the ones with ample data and immutable rules. The resulting cost savings can lead to it being used for more critical but less well-suited problems, such as forecasting and providing policy advice in times of extreme stress based on very limited or even irrelevant historical data. The risk is that AI becomes both irreplaceable and a source of systemic risk.
The third channel is AI misalignment, that is, difficulties in aligning the behaviour of AI engines with the objectives of their owners and the authorities. AI can lead to more market manipulation as it is likely better at identifying successful strategies and, beinmg incentivised by high-level, abstract objectives such as profit maximisation, not concerned with the ethical and legal consequences of its actions unless explicitly instructed. Even when not intentionally misbehaving, the superior performance of AI can destabilise the system because private sector AI excels in finding new and better ways of operating. While improved trading and risk management strategies will be individually profitable, they can be socially damaging. In times of extreme stress, AI decision-making can be much quicker and lead to undesirable behaviour, such as flights to safety, fire sales and investor runs.
Fourth, given the large startup costs involved in training models, AI is an increasing returns to scale technology. Human AI talent and computing resources are in limited supply, and crucial data for financial applications is often proprietary. All of these push the AI industry towards an oligopolistic market structure dominated by a few large vendors. This can lead to a monoculture in data analytics, inducing similar beliefs and actions among market participants and driving procyclicality, booms, and busts. It is a concern we have not appreciated the potential for AI-technology-data oligopolies in the recent wave of data vendor mergers.
The final channel of instability is AI evading control, which can happen because the financial system is nearly infinitely complex, making it very difficult for risk managers and the authorities to patrol. AI can help the authorities to monitor and keep the system stable but it also aids the forces of instability. The second factor is likely to dominate because an AI attempting to evade control only has to find one loophole to misbehave. At the same time, the risk controllers have to not only find all the weak points but also to monitor how AI interacts with each of them. As the private sector is better resourced than the public sector, and since those seeking to destabilise the system for profit or more nefarious purposes have an easier computational problem than the authorities, more powerful AI is likely to make the regulators’ job more difficult.
@@
Along the way, the use of AI in both the public and private sectors will need to be regulated. The reliance of both the regulated and regulators on the same AI vendors, along with challenges in data sovereignty and attribution of mistakes, demands special attention from the architects of future AI regulations. Meanwhile, the public sector’s use of AI in regulations raises tough questions. Who is accountable when a regulator’s AI makes decisions or provides crucial inputs for human decisions, and how can a regulated entity challenge decisions? The regulatory AI may not be able to explain its decision in terms that are intelligible to the subjects of regulatory actions, the AI’s human owner or to the judges involved in arbitration processes. To ensure themselves of the quality and safety of AI performance, authorities might find benchmarking AI against defined tasks a particularly fruitful way forward.
@@
The organisation of the paper is as follows. After the introduction, we present the conceptual challenges in Section 2, and follow that by a discussion on artificial intelligence and machine learning in Section 3. Section 4 is focused on the use of AI in the financial system, while Section 5 contains the five channels for how AI can destabilise the financial system. We then apply those to regulation of AI in Section 6. Section 7 concludes.
2 Conceptual challenges
@@
The use of AI in financial regulations is frustrated by the strength and weaknesses of AI, the complexity of the problem, and the particular incentives inherent in the financial system. Highly resourced financial market participants operate in very uncertain social environments that are subject to frequent structural changes, where neither the “rules of the game” are clearly spelled out nor the actors’ objectives known to each other. Sometimes, the participants can even change the rules to their advantage in a way that others only partially observe. Agents interact by competing against each other, cooperating with each other or even colluding to cheat or deceive others. Agents learn from interactions with the system changing their beliefs and hence how they will react in similar situations in the future. An AI that operates in such a system will encounter five conceptual challenges that can frustate the mission of the financial authorities seeking to use or regulate this technology.
@@ 2.1 Usefulness of data
Data should play to AI’s advantage as the financial system generates many petabytes of it daily. Every transaction is recorded, all decisions are documented, decision makers are monitored and recorded, and we can track processes over their lifetime. Financial institutions must report some of this data to the financial authorities, and the authorities can demand almost all of it later. One might expect it to be easy for AI to study the financial system in detail and identify all the causal relationships. That is true for most microprudential problems, but not macroprudential.
Start with the basic measurement. The standards for recording financial data are inconsistent, so different stakeholders might not record the same activity in the same way, leading to complex matching problems. Identification coding and database design can differ significantly. Financial institutions have a lot of legacy systems that were not set up with data collection and sharing in mind, rendering data collection, especially in a format that is standard across the industry and necessary for the authorities, costly and error prone. Fortunately, while real today, these problems are rapidly being overcome, not the least with the help of AI.
A bigger challenge is all the silos in the regulatory structure that hinder data sharing. Most data stays within a financial institution and is not shared. Even when shared, the financial institution might retain copyright, allowing it to control who sees the data so that it might be available for compliance but not for broader objectives such as financial stability. Furthermore, financial system data are collected by authority silos where data sharing is limited. There might be restrictions on data sharing between the supervisory and statistical unit of a central bank, between authorities in the same country or between jurisdictions. These problems were made clear in the crisis in 2008, where nobody had an overview of the aggregate market for structured credit. The situation has improved somewhat since then due to mandatory trade reporting for many derivatives transactions and increasing availability of data on cash and repo transactions for bond markets. However, silos persists and digesting the large amount of data remains a challenge.
Finally, when it comes to the most serious events, systemic financial crisis, the events under consideration are, fortunately not frequent. The typical OECD country only suffers a systemic crisis one year out of 43, according to the crisis database maintained by Laeven and Valencia (2018).
These three issues, data quality, silos and rarity of events, are usually not all that important for microprudential regulations. Not so for macropru where they amplify each other. That, in turn, impacts both the design of macroprudential regulations, enforcement, and crisis resolution.
@@ 2.2 Unknown-unknowns
The second conceptual problem arises from most crises being unknown-unknown events that are both unique and infrequent. It is almost axiomatic that the type of event a macroprudential authority is concerned with plays out in unexpected ways, otherwise precautionary actions would have been taken to avoid a crisis.
Unfortunately, from the point of view of AI learning from these crises, the details of a given crisis are, in important aspects, unique to it. Fundamentally, every crisis is caused by the same set of fundamental vulnerabilities, all of which act as crisis amplifiers. Financial institutions that use high degrees of leverage that render them vulnerable to shocks, self preservation in times of stress leading market participants to prefer the most liquid assets and system opacity and complexity and asymmetric information causing market participants to mistrust each other in times of heightened uncertainty. However, these vulnerabilities are essentially conceptual, and when it comes to designing regulations to prevent stress and mitigating it when it happens, the authorities have to focus on details. Those details are unique to each crisis. That is almost self evident because the supervisors would have prevented a crisis if they were not.
While the authorities can scan the system for specifics that cause vulnerabilities, their job is frustrated by the almost infinite complexity of the financial system so that the supervisors can only patrol a small part of it. Even if supervisors, AI or human, could monitor all threat scenarios and assign a probability to each — an impossible task — they still have the problem of picking notification thresholds. The system’s complexity and measurement noise mean that the number of notifications would be very large, with mostly false positives. Furthermore, such intrusive monitoring might sharply curtail desirable risk taking because of the false positives, and be seen as socially unacceptable.
This uniqueness of crises creates particular problems for the designers of macroprudential regulations because they really only know what data is useful after an event. That was, for example, the case in the crisis in 2008, where we only realised afterwards that sub-prime mortgages being put into structured credit products — where the banks held onto the most senior and junior tranches where risk modelling was extremely poor — was the key channel for the crisis. Obvious after the event but practically impossible to discover before. When the analyst has an almost infinite amount of signals and an enormous amount of false positives, it is very difficult, to the point of impossible, to identify which data is useful until after a crisis event is already underway. It is too late to have preventative regulations in place by this time.

### 03 Feb 2024 v4

@@ 1 Introduction
Artificial intelligence (AI) is transforming the financial system, promising improved efficiency, robustness and impartiality at much lower costs than existing arrangements. AI also threatens financial stability when AI vulnerabilities viciously interact with economic fragilities.
@@
While there is no single notion of what AI is, it is helpful to see it as a computer algorithm performing tasks usually done by humans. AI differs from machine learning (ML) and traditional statistics in that it not only provides quantitative analysis but also gives recommendations and makes decisions. [^1] For general overview of AI Russel (2019) is particularly useful. For a more technical review of the underlying ML techniques, see Murphy (2023) and Prince (2023). Today’s AI excels at extracting patterns from large, unstructured datasets for use in classification and prediction tasks. But they do not reason in the same way humans do. Once AI achieves proficiency at a given task — has been trained — it can advise human decision makers or make decisions autonomously. Norvig and Russell (2021) list a number of possible definitions of AI. Among these, AI as a rational maximising agent resonates with the economic notion of utility maximising agents and, hence, is particularly helpful in the analysis of AI use in the financial system.
A financial authority entrusted with regulating the financial system has two objectives. The first is microprudential regulation, or micropru, concerned with day-to-day issues such as risk management, consumer protection and fraud. AI is generally beneficial for micropru as data is ample, the rules are mostly fixed on the timescale decisions are made and the cost of mistakes is small.
The second policy objective is macroprudential regulations, or macropru, focusing on broad-picture issues relating to financial stability. The emphasis here is decidedly long run, both the avoidance of large losses and systemic financial crises years and decades into the future, as well as the resolution of crises when they occur. Macropru is more difficult to execute and less accurate than micropru, as events to be controlled are infrequent and often unique. That implies considerable uncertainty about the ability of AI and the dangers arising from its use, as noted by Danielsson, Macrae and Uthemann (2022).
It is helpful to think of AI’s application to financial policy on a spectrum. At one end, we have a problem with ample data and clear and simple objectives, playing to the strength of AI and it’s learning algorithms. As the frequency of events drops and rules are increasingly mutable, the AI advantages erode until we reach once-in-a-working-lifetime events, such as the resolution of a systemic financial crisis, where AI, though very valuable, can threaten the system’s stability.
@@
The use of AI in the financial system is growing rapidly. Private sector financial institutions apply it to tasks such as risk management, credit allocation, fraud detection and regulatory compliance. Even if many traditional banks have been reticent in its adoption, some employ large AI teams, and challenger banks have been particularly enthusiastic about its use. Significant cost savings and efficiency improvements in a highly competitive landscape will likely rapidly drive the further adoption of AI in the private sector.
@@
The financial authorities need to respond, and most are formulating AI policy (see, e.g. Kiarelly et al., 2024; Moufakkir, 2023). While some authorities might prefer a slow, deliberative and conservative approach to AI, that will not be tenable for three reasons. First, an authority that persists in regulating with traditional methods may find itself outmanoeuvred. Second, a policy of not using AI for high-level decisions will likely be undermined by AI being adopted by stealth. Finally, AI may be considered essential for advising on a critical task. Even if an authority wanted to limit AI to basic advisory roles, there might not be much difference between AI making decisions and AI providing crucial advice, particularly if its internal representation of the system is not intelligible to its human operators.
@@
The usefulness of AI for the financial authorities is directly affected by five conceptual economic challenges that can viciously interact with AI vulnerabilities. The first, perhaps paradoxically, is data. After all, the financial system may appear to be the ideal use case for AI because it creates vast amounts of data, leaving plenty for AI to train on. However, financial data collection is incomplete, and recorded data are often inconsistently or even inaccurately measured, leading to non-representative samples. What is more, financial crises are rare, meaning there is little data available on the most important events for macropru.
@@
The second challenge arises from the uniqueness of crises. While major financial crises share common fundamental features, every crisis is unique because the driving factors are specific to the institutional structure and the particular regulatory and political regimes in place. Rarity and uniqueness in a sparsely monitored system imply that systemic crises can be seen as “uncertain” in Frank Knight’s (1921) classification and thus do not fit well into the conventional ML paradigm.
@@
AI’s third conceptual challenge relates to how the financial system responds to control, echoing Goodhart’s (1974) law and the Lucas (1976) critique. New regulations change the structure of the financial system in ways that are hard to predict, making it difficult for AI to learn from historical data about the possible consequences of regulatory actions.
@@
The fourth conceptual problem stems from a lack of clarity on what objectives the AI should pursue. This is particularly problematic for macropru, where the regulators’ problem is often described in abstract high-level terms such as “ensure financial stability”. The objectives of the most important macropru actions only emerge in times of crisis, in a process led by the political leadership.
The final conceptual challenge relates to difficulties in aligning AI’s incentives with the authorities’ objectives. The purpose of financial regulations is to align the private sector’s incentives with society at large. The interactions between financial supervisors and private sector decision-makers can be seen as a difficult principal-agent (PA) relationship because of incomplete contracting over a latent variable, risk, in a world with explicit and implicit government guarantees. With AI, the one-sided financial institution-regulator problem becomes two-sided, amplifying the already significant existing PA problem in financial regulation.
@@
The literature on AI has identified particular risks that AI engines pose; see, for example, Barnichon et al. (2022), and several researchers, such as Weidinger et al. (2022), Bengio et al. (2023) and Shevlane et al. (2023) identify societal risks arising from AI. When studying how these interact with the economic conceptual challenges listed above, we find four channels for how AI use may destabilise the financial system.
@@
The first channel is the malicious use of AI by its human operators, a particular concern in the financial system because it is replete with highly resourced profit-maximising economic agents not too concerned about the social consequences of their activities. Such agents can bypass controls and change the system in a way that benefits them while being difficult for competitors and regulators to detect. They may even deliberately create market stress, which is highly profitable for those forewarned. These agents either directly manipulate AI engines or use them to find loopholes to evade control. Both are easy in a financial system that is effectively infinitely complex. Such activities can be socially undesirable and even be against the interests of the institution employing the operator of the AI engine. We expect the most common malicious use of AI will be by employees of financial institutions, careful to stay on the right side of the law. AI will likely also facilitate illegal activities, such as rogue traders and criminals, as well as terrorists and nation-states aiming to create social disorder.
The second channel is due to the users of AI being both misinformed about its abilities and strongly dependent on it. That is most likely when data-driven algorithms, such as those used by AI, are asked to extrapolate to areas where data is scarce and objectives unclear, which is very common in the financial system. AI engines are designed to provide advice even when they have very low confidence about the accuracy of their answer. They can even make up facts or present arguments that sound plausible but would be considered flawed or incorrect by an expert, both instances of the broader phenomenon of AI hallucination. The risk is that AI will present confident recommendations about outcomes they know little about. To overcome that, the engines will have to provide an assessment of the statistical accuracy of their recommendations. Here, it will be helpful if the authorities overcome their frequent reluctance to adopt consistent quantitative frameworks for measuring and reporting on the statistical accuracy of their data-based inputs and outputs.
The third channel emerges from the difficulties in aligning the objectives of AI with those of its human operators. While we can instruct AI to follow our instructions, there is no guarantee it will actually do so. It is impossible to pre-specify all the objectives AI has to meet, a problem since AI is very good at manipulating markets and being incentivised by high-level objectives such as profit maximisation is not concerned with the ethical and legal consequences of its actions unless explicitly instructed. The superior performance of AI can destabilise the system even when it is only doing what it is supposed to do. That is particularly problematic in times of extreme stress when the objective of financial institutions, and hence the AI working for them, is survival, amplifying existing destabilising behaviour, such as flights to safety, fire sales and investor runs. More generally, AI will find it easy to evade oversight because it is very difficult to patrol a nearly infinitely complex financial system. The authorities have to contend with two opposing forces. AI will be very helpful in keeping the system stable but, at the same time, aids the forces of instability. We suspect the second factor dominates. The reason is that AI attempting to evade control only has to find one loophole to misbehave, while the supervisors not only need to find all the weak points but also monitor how AI interacts with each of them and then effectively implement corrective measures. That is a very difficult computational task, made worse by the private sector having access to better computational resources than the authorities. The more we use AI, the more difficult the computational problem for the authorities becomes.
The final channel emanates from the business models of those companies designing and running AI engines exhibits increasing returns to scale. AI analytics businesses depend on three scarce resources: compute, human capital and data, pushing the AI industry towards an oligopolistic market structure dominated by a few large vendors. The end result is amplified procyclicality and more booms and busts. Suppose the authorities also depend on the same AI engine for its analytics, which seems likely. In that case, they may not be able to identify the resulting fragilities until it is too late. In other words, the oligopolistic nature of the AI analytic business increases systemic financial risk. It is a concern that neither the competition nor the financial authorities appear to have fully appreciated the potential for increased systemic risk due to oligopolistic AI technology in the recent wave of data vendor mergers.
@@
Along the way, the use of AI in both the public and private sectors will need to be regulated. The reliance of both the regulated and regulators on the same AI vendors, along with challenges in data sovereignty and attribution of mistakes, demands special attention from the architects of future AI regulations. Meanwhile, the public sector’s use of AI in regulations raises tough questions. Who is accountable when a regulator’s AI makes decisions or provides crucial inputs for human decisions, and how can a regulated entity challenge decisions? The regulatory AI may not be able to explain its decision in terms that are intelligible to the subjects of regulatory actions, the AI’s human owner or the judges involved in arbitration processes. To ensure the quality and safety of AI performance, authorities might find benchmarking AI against defined tasks a particularly fruitful way forward.
We do not want to overemphasise these issues. We suspect that the benefit of AI will be overwhelmingly positive in the financial system. However, the authorities must be alive to these threats and adapt regulations to meet them. The ultimate risk is that AI becomes both irreplaceable and a source of systemic risk before the authorities have formulated the appropriate response.
@@
The organisation of the paper is as follows. After the introduction, we present the conceptual challenges in Section 2 and follow that with a discussion on artificial intelligence and machine learning in Section 3. Section 4 is focused on the use of AI in the financial system, while Section 5 contains the five channels for how AI can destabilise the financial system. We then apply those to the regulation of AI in Section 6. Section 7 concludes.
2 Economic conceptual challenges
@@
Any attempt at controlling the financial system is frustrated by the presence of highly resourced financial market participants operating in very uncertain social environments that are subject to frequent structural changes. The rules of the game are usually unclear, and the actors often do not know each other’s objectives. The participants can even change the rules to their advantage in a way that others only partially observe. Agents interact by competing, cooperating or even colluding with each other, cheating and deceiving along the way. They learn from interactions with the system and are continually changing their beliefs and, hence, how they will react in similar situations in the future.
Any authority, be it AI or human, government regulators or internal controllers will have to content with five economic conceptual challenges that get in the way of their mission. However, the inherent characteristics of AI make AI controllers particularly affected by these challenges.
@@ 2.1 Usefulness of data
Data should play to AI’s advantage as the financial system generates many petabytes of it daily. Every transaction is recorded, all decisions are documented, decision makers are monitored and recorded, and we can track processes over their lifetime. Financial institutions must report some of this data to the financial authorities, and the authorities can demand almost all of it later. One might expect this ocean of data would make it easy for AI to study the financial system in detail and identify all the causal relationships. That is true for most microprudential problems, but not the macroprudential.
Start with the basic measurement. The standards for recording financial data are inconsistent, so different stakeholders might not record the same activity in the same way, leading to complex matching problems. Identification coding and database design can differ significantly. Financial institutions have a lot of legacy systems not set up with data collection and sharing in mind, rendering data collection, especially in a format that is standard across the industry and necessary for the authorities, costly and error prone. Fortunately, while real today, these problems are rapidly being overcome, not the least with the help of AI.
A bigger challenge is all the silos in the regulatory structure that hinder data sharing. Most data stays within a financial institution and is not shared. Even when shared, the financial institution might retain copyright, allowing it to control who sees the data so that it might be available for compliance but not for broader objectives such as financial stability. Furthermore, financial system data are collected by authority silos where data sharing is limited. There might be restrictions on data sharing between the supervisory and statistical unit of a central bank, between authorities in the same country or between jurisdictions. These problems were made clear in the crisis in 2008, where nobody had an overview of the aggregate market for structured credit. The situation has improved somewhat since then due to mandatory trade reporting for many derivatives transactions and the increasing availability of data on cash and repo transactions for bond markets. However, silos persist and digesting the large amount of data remains a challenge.
Finally, when it comes to the most serious events, systemic financial crisis, the events under consideration are, fortunately not frequent. The typical OECD country only suffers a systemic crisis one year out of 43, according to the crisis database maintained by Laeven and Valencia (2018). However, that fortunate low frequency of crises frustrates the data driven analytics AI depends on.
These three issues, data quality, silos and rarity of events, are usually not all that important for microprudential regulations. Not so for macropru, where they amplify each other, negatively impacting the design of macroprudential regulations, enforcement, and crisis resolution.
@@ 2.2 Unknown-unknowns
The second conceptual problem arises from most crises being unknown-unknown events that are both unique and infrequent. It is almost axiomatic that the type of event a macroprudential authority is concerned with plays out in unexpected ways. Otherwise, precautionary actions would have been taken to avoid a crisis.
Unfortunately, from the point of view of AI learning from these crises, the details of a given crisis are, in important aspects, unique to it. Fundamentally, every crisis is caused by the same set of fundamental vulnerabilities, all of which act as crisis amplifiers. Financial institutions that use high degrees of leverage that render them vulnerable to shocks, self-preservation in times of stress, leading market participants to prefer the most liquid assets, and system opacity, complexity and asymmetric information, causing market participants to mistrust each other in times of heightened uncertainty. However, these vulnerabilities are essentially conceptual, and when it comes to designing regulations to prevent stress and mitigating it when it happens, the authorities have to focus on details. Those details are unique to each crisis. That is almost self evident because the supervisors would have prevented a crisis if they were not.
While the authorities can scan the system for specific causes of vulnerabilities, their job is frustrated by the almost infinite complexity of the financial system. The supervisors can only patrol a small part of that infinitely complex system. Even if the supervisors, AI or human, could monitor all threat scenarios and assign a probability to each — an impossible task — they still have the problem of picking notification thresholds. The system’s complexity and measurement noise means that the number of notifications would be very large, with mostly false positives. Furthermore, such intrusive monitoring might sharply curtail desirable risk taking because of the false positives, and be seen as socially unacceptable.
This uniqueness of crises creates particular problems for the designers of macroprudential regulations because they generally only learn what data is most useful after a stress event. That was, for example, the case in the crisis in 2008, where we only realised afterwards that sub-prime mortgages being put into structured credit products — the banks held onto the most senior and junior tranches where risk modelling was extremely poor — was the key channel for the crisis. Obvious after the event but practically impossible to discover before. When the analyst faces an almost infinite number of signals and an enormous number of false positives, it is very difficult, to the point of impossible, to identify which data is useful until after a crisis event is already underway. It is too late to have preventative regulations in place by this time.